OpenStudy (anonymous):

x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values

OpenStudy (anonymous):

heres the rest of the question... Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received.

ganeshie8 (ganeshie8):

\[\large \rm x - 62,500 = 0.25y\] plugin \(\large \rm y = 125,000\) and solve \(\large \rm x\)

OpenStudy (anonymous):

thx!

ganeshie8 (ganeshie8):

are you sure you can work the rest ?

OpenStudy (anonymous):

no not really its hard can you help me..

ganeshie8 (ganeshie8):

\[\large \rm x - 62,500 = 0.25y\] you're given a total of 125,000 ppl have voted so plugin y = 125,000 : \[\large \rm x - 62,500 = 0.25(125,000)\]

ganeshie8 (ganeshie8):

do u have calculator ?

OpenStudy (anonymous):

oh no my bad..